Duty Calls has developed a project that requires an immediate investment of $210. The project's long-term cash flows are $80 for 4 consecutive years beginning in one year, and the project has a required rate of return of 8%. Based on these estimates, what is the project's NPV.

Project NPV: $

Place your answer in dollars and cents. Do NOT include a dollar sign or a comma in your NPV. For example, an answer of ten thousand should be placed as 10000. If applicable, indicate negative amounts with a minus sign in front of the number.

Work your analysis using at least 4 decimal places of accuracy.
Answer #1

fr 54.9701 G358 B C D 349 350 351 352 353 354 Year Cashflows Pv@8% $ (210.00) 80.00 80.00 80.00 $ 80.00 IS PVCF 1 $ 0.9259259for 54.9701 G358 B 349 350 Year Pv@8% Cashflows -210 352 353 354 355 =F352/1.08 =F353/1.08 =F354/1.08 =F355/1.08 PVCF =E352*F